Circular No. 6781

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997ea IN AN ANONYMOUS GALAXY The University of Texas McDonald Observatory Supernova Search

Team (M. T. Adams, T. Montemayor, D. A. Howell, J. C. Wheeler, M. H. Ward, and W. Wren) report their discovery of a supernova (R ∼ 19) on three Kron-Cousins R-band CCD images of Abell 595 (z = 0.0689) taken on Nov. 30.39–30.40 UT with the 0.76-m telescope. SN 1997ea is located at α = 7^{h}48^{m}37^{s}.70, δ = +52^{o}13′22′′.8 (equinox 2000.0), which is ∼ 9′′ east and 3′′ north of the nucleus of the apparent host elliptical galaxy. A finder

chart is available at http://draco.as.utexas.edu/ward/sn2/sn.html, or

by anonymous ftp from draco.as.utexas.edu. BVRI photometry of SN 1997ea was acquired on Dec. 3.22–3.45; it is not visible on R-band images (limiting R ∼ 21) acquired on Oct. 30.41–30.43.

P. Garnavich, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that spectra of SN 1997ea were obtained with the Multiple Mirror Telescope on Dec. 5.4 UT and show that it is a type-Ia supernova, ∼ 2 weeks past maximum. The recession velocity estimated from the broad features in the spectrum is ∼ 19 500 km/s, consistent with the redshift of Abell 595.

SUNGRAZING COMETS C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports that S. Stezelberger has discovered two probable Kreutz sungrazing comets, each of mag 5 and showing a tail, in C3 and C2 coronagraphic data. Measurements by D. A. Biesecker and St. Cyr have been reduced by G. V. Williams and are published in detail on MPEC 1997-X08 and -X09, together with orbit solutions by B. G. Marsden.

1997 UT  α_{2000}  δ_{2000}  Comet
Nov. 21.725  15^{h}33.^{m}8  −25^{o}00′  C/1997 W1
Dec. 1.665  16 26.0  −26 27  C/1997 X1

BW CIRCINI R. Soria, M. S. Bessell, and P. Wood, Mount Stromlo Observatory

report that J- and K-band infrared images of BW Cir, counterpart of the blackhole candidate GS 1354–64 (IAUC 6772, 6774, 6775, 6779), were obtained with the Australian National University 2.3-m telescope at Siding Spring Observatory on Nov. 20.73 UT by P. Wood, yielding the following preliminary magnitudes: J = 15.35 ± 0.05; K = 13.95 ± 0.05.

1997 December 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green